PID based parameters given in handwritten imageKn=0.5+ 0.07 0.57 mA Vt = 0.6 to.02 = 0.62 V R 10. (1.1)= 11kR V 4.17 V V = 4.7 V R R V20 o V₃ M1 M2 V find 100 MA (b) Redraw
Answer #1

Que Griven data kn = 0.57 mA /v2 Vt = 0.6+0.02= 0.62 V, = 4.7 R = 11 kn Given Craut diagram Vi = 4.7 V, = 4.7V R = lika R} =0-09615 - (unsi 0.62)? Vasi 0.62 0-31 Versi 0-9 3 Volt So vas, - vaso =0.93V and 3 4. 7-11 X 50 X 10 كما V4 = كما 4-15 Volt V